# LIFT-GS: Cross-Scene Render-Supervised Distillation for 3D Language Grounding

### v1

@@ Abstract
Our approach to training 3D vision-language understanding models is to train a feedforward model that makes predictions in 3D, but never requires 3D labels and is supervised only in 2D, using 2D losses and differentiable rendering. The approach is new for vision-language understanding. By treating the reconstruction as a “latent variable”, we can render the outputs without placing unnecessary constraints on the network architecture (e.g. can be used with decoder-only models). For training, only need images and camera pose, and 2D labels. We show that we can even remove the need for 2D labels by using pseudo-labels from pretrained 2D models. We demonstrate this to pretrain a network, and we finetune it for 3D vision-language understanding tasks. We show this approach outperforms baselines/sota for 3D vision-language grounding, and also outperforms other 3D pretraining techniques. Project page: https://liftgs.github.io.
@@ 1 Introduction
When a user mentions the keys by the door or the blue mug on the table, they use language to indicate a specific set of objects and 3D locations in space. Such 3D language grounding provides a particularly natural interface for people to communicate about their surroundings.
@@
Figure 1: LIFT-GS. We train a 3D vision language grounding model (3D VLG) with point clouds and language inputs by distilling from 2D foundation models without any 3D supervisions.
For AI systems operating in physical spaces, identifying the set of 3D masks or bounding boxes indexed by a language query represents a core functionality; with applications across autonomous navigation, robotic manipulation, and AR/VR. Yet, despite its importance and recent advances in both 2D vision-language grounding and 3D reconstruction, the gap between human and machine performance in 3D Vision-Language Grounding remains one of the most significant challenges in embodied AI.
3D vision-language grounding (3D VLG) is severely constrained by data scarcity. Generative multimodal models are routinely trained on billions to trillions of tokens from language (Achiam et al., 2023; Touvron et al., 2023), images (Radford et al., 2021; Labs, 2023), videos (Polyak et al., 2024; Brooks et al., 2024). Nearer to the task of 3D VLG, models trained for promptable image and video mask segmentation use millions to billions of labeled masks (Kirillov et al., 2023; Ravi et al., 2024). In contrast, existing 3D VLG models are limited to training on only tens of thousands of labeled 3D masks and scenes, restricting their ability.
In this paper, we propose a scalable pretraining approach for 3D VLG that enables grounding language into a set of target 3D masks; without requiring 3D mask labels. Language-Indexed Field Transfer with Gaussian Splatting (LIFT-GS), is supervised on 2D frames directly. It uses differentiable rendering to render the predicted 3D features and masks into 2D frames (e.g. via Gaussian Splatting (Kerbl et al., 2023)). LIFT-GS is then distilled directly from powerful frame-based foundation models, using pseudolabels from SAM (Kirillov et al., 2023), CLIP (Radford et al., 2021), and Llama 3.2 (Meta AI, 2024).
LIFT-GS conducts both Promptable 3D Segmentation and Reconstruction. The render-supervised formulation is highly flexible; as it imposes essentially no constraints on the model input or network architecture. Therefore, in order to demonstrate the effectiveness of differentiable rendering for structured prediction tasks, we choose model inputs and meta-architectures that are scalable and offer advantages for 3D vision-language grounding.
As visual input, LIFT-GS exclusively utilizes point clouds (positions and colors) to align with the preferences for many modern embodied AI applications that operate in real time. Point clouds from SLAM, for example, serve as widely adopted input in robotics and AR/VR. By relying on raw point clouds rather than 2D feature-enhanced point clouds like ConceptFusion (Jatavallabhula et al., 2023b), LIFT-GS eliminates the need for preprocessing and feature fusion during inference, reducing inference time from approximately 60 seconds per frame to just 1 second for the forward pass.
LIFT-GS follows the common MaskFormer (Cheng et al., 2021b) meta-architecture. However, the output tokens from the Encoder and a Mask Decoder are instead interpreted as 3D Gaussians rather than image patches. The entire pipeline is then trained end-to-end by optimizing the 2D loss between the rendered 2D masks and the pseudo-grounding masks.
We validate the effectiveness of LIFT-GS on two popular 3D vision-language grounding (3D VLG) downstream tasks: open-vocabulary 3d instance segmentation and 3D referential grounding. The results show that distillation provides significant improvements over models trained from scratch, and achieves state-of-the-art performance on both tasks. More importantly, the approach consistently improves from both more data and better 2D foundation models, which indicates the potential for further improvement.
Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data. This somewhat counterintuitive observation indeed matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021), where pretraining shifts the power-law scaling in log-dataset size to the left by that constant factor. The fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, underscores that 3D VLG models are currently operating in the data scarce regime.
The seeming universality of scaling laws across modalities, and the effectiveness of LIFT-GS pretraining on in 3D VLG settings, suggests that the underlying tools may also be useful at least in other 3D understanding settings. Specifically, render supervised framework can be used with essentially any 3D/4D task or model, provided the results are renderable. This can be combined with knowledge distillation from frame-based models in order to overcome data scarcity in 3D, and LIFT-GS specifically addresses a scarcity of 3D grounded masks. For summary, our contributions are:
  • Differentiable rendering as a tool for training large-scale 3D promptable segmentation models. Instead of proposing a new architecture, the contribution is to train the model using differentiable rendering with a structured 2D grounding loss.
  • A pseudo-labeling strategy for distilling a 2D vision-language grounding pipeline into 3D versions. The pretrained 2D models are needed only for pseudo-labeling and they are not needed during inference.
  • State-of-the-art performance and realistic evaluations. We demonstrate the effectiveness of the approach using sensor pointclouds, common in embodied settings. Rigorous experiments show state-of-the-art performance and reveal scaling properties.
Figure 2: Illustration of 3D Referential Grounding Task. Given an object description, the task requires the model to predict 3D masks for all mentioned noun phrases while ensuring that the located objects satisfy the semantic relationships described in the entire sentence.
@@ 2 Related Work
3D Vision-Language Grounding 3D Vision-Language Grounding (3D VLG) refers to the task of mapping language descriptions of objects in a scene, to a set of corresponding 3D masks or bounding boxes that indicate the location of the objects in the observed scene. Despite its importance, there is very little annotated data for 3D VLG available — mainly because there are very little 3D data with annotated masks or bounding boxes period, as well as language annotation for each instance. This provides a major challenge for essentiall all existing models, since they require access to use ground-truth 3D masks or bounding boxes for trainings (Yuan et al., 2021; Roh et al., 2021; Yang et al., 2021; Zhu et al., 2023c, 2024), and many require them during inference in the form of proposal masks or bounding boxes (Fang et al., 2024; Zhang et al., 2023b). This lack of annotated data is a major challenge for all existing models, as the number of 3D scenes with annotated masks or 3D bounding boxes is in the thousands (Dai et al., 2017; Yeshwanth et al., 2023; Somasundaram et al., 2023), compared to the millions of images and videos used for image and video segmentation (Kirillov et al., 2023).
Unlike prior approaches, LIFT-GS introduces a render-supervised approach for 3D VLG. It uses differentiable rendering to train directly on image masks, without requiring 3D bounding boxes or masks during training or inference. This formulation is architecture agnostic and allows the model to be trained using only image and language losses, without the need for 3D masks or bounding boxes. LIFT-GS demonstrates this by training an end-to-end 3D feedforward model. LIFT-GS takes as input a RGB pointcloud and a language utterance as input, and outputs 3D masks.
3D Instance Segmentation 3D instance segmentation refers to the task of predicting masks for a set of instance in a 3D scene, not necessarily given a language prompt. For example, using a predefined set of category labels (Qian et al., 2021; Schult et al., 2022; Hou et al., 2018) or interactive segmentation using input “prompts” (Osep et al., 2024; Zhou et al., 2024; Ma et al., 2024; Chen et al., 2024a).
Because of the lack of 3D data with masks with which to train a 3D segmentation model, a class of recent work has emerged that trains no predictive 3D segmentation model at all, but instead directly lifts powerful pretrained image and video models to 3D using per-scene optimization and multiview geometry. For example, by using depth-unprojection with heuristic view-merging strategies such as voxel-voting (Jatavallabhula et al., 2023a; Zhou et al., 2024; Xu et al., 2023b). Another more recent approach to lift 2D segmentation to 3D is by using differentiable rendering to optimize the 3D representations to minimize the 2D loss between rendered outputs and ground-truth RGB values, 2D features (Kerr et al., 2023; Kim et al., 2024; Dou et al., 2024; Chen et al., 2024b; Gu et al., 2024), or masks (Cen et al., 2023; Xu et al., 2023a). These approaches leverage the large-scale data available for image and video segmentation, but use a fixed per-scene lifting process that does not improve with more training data and takes a long time to run (e.g. minutes per scene on a H100).
These lifted 3D pseudolabels can be used to partially address the lack of 3D data with ground-truth masks, by pre-training a 3D instance segmentation model on the 3D pseudolabels (Genova et al., 2021; Peng et al., 2023). However, this fixed lifting pipeline introduces accumulated errors from inaccuracies in 3D reconstruction and label-merging strategies, creating a bottleneck for overall performance.
LIFT-GS instead trains a promptable 3D segmentation model using only 2D losses, eliminating the need for a fixed lifting process and label-merging strategies. Not relying on heuristic lifting methods or specialized losses (e.g., cross-mask contrastive loss), LIFT-GS is trained on a large-scale dataset using the same mask losses (Dice, Cross-Entropy Focal) as the original frame-level segmentation models. As shown in Appendix Section A.2, LIFT-GS features trained directly on 2D pseudolabels outperform these lifted 3D pseudolabels. And we show effective data scaling in Section 4.4.
Render-Supervised 3D While photometric losses using differentiable rendering (Mildenhall et al., 2020; Kerbl et al., 2023) were originally used for per-scene optimization, render-supervision is emerging as a powerful and generic technique to train predictive 3D models using only 2D losses. For 3D reconstruction, 2D photometric losses can be used to train strong feedforward models (Hong et al., 2024; Tang et al., 2024). PonderV2 (Zhu et al., 2023a) instead uses render-supervision for representation learning, by adding additional constrastive losses on the rendered pixels, using ground-truth category labels to pretrain an Encoder.
LIFT-GS introduces render-supervision for 3D VLG, using mask- and grounding-losses typically used for 2D VLG. LIFT-GS jointly trains an Encoder and Mask Decoder, and for the first time demonstrates using differentiable rendering to train a 3D mask decoder.
@@ 3 Method
LIFT-GS is a (pre)training pipeline for 3D vision-language grounding models that uses only 2D supervision. It accomplishes this using differentiable rendering and structured 2D losses, and uses pseudo-labeling for frame annotations.
@@ 3.1 Task Formulation
The formulation of 3D Vision-Language Grounding used in LIFT-GS is shown in Figure 2, and based on MDETR (Kamath et al., 2021). The model takes as input a language query, such as “the black chair close to the table near the wall and a pointcloud input. The objective is to predict 3D gaussian masks, for all mentioned noun phrases (chair, table, wall); ensuring that the predicted masks adhere to the semantic relationships described in the language query, via a mask-to-query correspondence matrix.
As shown in Figure 4, LIFT-GS takes as input a point cloud, $\mathbf{P}$, and language query embeddings, $\mathbf{Q}$. It outputs 3D feature Gaussians $\mathbf{G}$ with a correspondence matrix $\mathbf{C}$.
@@
The pointcloud $\mathbf{P}\in\mathbb{R}^{|P|{\times}6}$ contains positions and RGB colors. LIFT-GS is trained with pointclouds from unprojected posed multi-view RGBD, but any pointcloud reconstruction will work; including pointmap-based methods that do not require depth or camera pose (Wang et al., 2023; Yang et al., 2025; Wang et al., 2025). This makes the method highly flexible, applicable across diverse scenarios, from sparse-view settings to extremely long video sequences.
The Language queries are encoded into embeddings $\mathbf{Q}\in\mathbb{R}^{|Q|{\times}F_{Q}}$, i.e. via CLIP (Radford et al., 2021) in our design. Each embedding can be either from a single word or a full sentence, depending on the desired ganulariry of grounding.
The model predicts 3D Gaussians $\mathbf{G}\in\mathbb{R}^{|G|{\times}|F_{G}|}$ with auxilliary features, enabling the differentiable rendering of 3D information into 2D. In LIFT-GS, the first part of channels of $F_{G}$ are standard Gaussian splatting parameters (i.e. positions $\mu$, covariances $\Sigma$, 3rd-degree SH coefficients $\psi$). The remaining channels are renderable auxiliary 3D attributes: predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|{\times}m}$, and renderable auxiliary features: predicted features $\mathbf{F}\in\mathbb{R}^{|G|{\times}F}$.
Following the trend in mask grounding (Liu et al., 2023), the predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|\times m}$ contains $m$ predicted masks for each Gaussian, where $m$ is the mask proposal number. Consequently, the rows in the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$ represent the log-probability that the $i$-th predicted mask corresponds to each embedding of $\mathbf{Q}$.
3.2 Training 3D with 2D Supervision
Predicting a 3D Gaussian Splatting with desired 3D attributes enables training a 3D model using only 2D supervision, eliminating the need for large-scale 3D annotations. In this section, we first discuss the 2D losses used for training and then describe the process of generating pseudo-labels when perfect labels are unavailable. We show ablations for both losses and pseudo-labels in the experiments.
LIFT-GS utilizes three 2D losses to train the model: $\mathcal{L}_{\text{RGB}}$, $\mathcal{L}_{\text{feat}}$ and $\mathcal{L}_{\text{ground}}$. The rendered image, feature map, and 2D mask are denoted in the equations below as $\tilde{I}\in\mathbb{R}^{H\times W\times 3}$, $\tilde{\mathbf{F}}_{\text{2D}}\in\mathbb{R}^{H\times W\times F}$, and $\tilde{\mathbf{M}}_{\text{2D}}\in\mathbb{R}^{H\times W\times m}$ , respectively. Their corresponding ground-truth counterparts are $I$, $\mathbf{F}_{\text{2D}}$, and $\mathbf{M}_{\text{2D}}$, where $\mathbf{M}_{\text{2D}}\in\mathbb{R}^{H\times W\times K}$.

### v2

# From Thousands to Billions: 3D Visual Language Grounding via Render-Supervised Distillation from 2D VLMs

@@ Abstract
3D vision-language grounding faces a fundamental data bottleneck: while 2D models train on billions of images, 3D models have access to only thousands of labeled scenes–a six-order-of-magnitude gap that severely limits performance. We introduce LIFT-GS, a practical distillation technique that overcomes this limitation by using differentiable rendering to bridge 3D and 2D supervision. LIFT-GS predicts 3D Gaussian representations from point clouds and uses them to render predicted language-conditioned 3D masks into 2D views, enabling supervision from 2D foundation models (SAM, CLIP, LLaMA) without requiring any 3D annotations. This render-supervised formulation enables end-to-end training of complete encoder-decoder architectures and is inherently model-agnostic. LIFT-GS achieves state-of-the-art results with 25.7% mAP on open-vocabulary instance segmentation (vs. 20.2% prior SOTA) and consistent 10-30% improvements on referential grounding tasks. Remarkably, pretraining effectively multiplies fine-tuning datasets by 2×, demonstrating strong scaling properties that suggest 3D VLG currently operates in a severely data-scarce regime. Project page: https://liftgs.github.io.
@@ 1 Introduction
When a user mentions the keys by the door or the blue mug on the table, they use language to indicate a specific set of objects and 3D locations in space. Such 3D language grounding provides a particularly natural interface for people to communicate about their surroundings. For AI systems operating in physical spaces, identifying the 3D masks or bounding boxes indexed by language queries represents a core functionality, with applications across autonomous navigation, robotic manipulation, and AR/VR.
@@
Figure 1: LIFT-GS  Overview. We train a powerful 3D vision language grounding model (i.e., 3D mask decoder) with point clouds and language as inputs by learning from 2D VLM foundation models without any 3D supervision.
Yet despite its importance, 3D vision-language grounding (3D VLG) faces a fundamental bottleneck: data scarcity. While 2D vision-language models are trained on billions of labeled images and masks (Achiam et al., 2023; Touvron et al., 2023; Radford et al., 2021; Labs, 2023), existing 3D VLG models have access to only thousands of labeled 3D scenes and masks. This six-order-of-magnitude gap in data availability severely limits the capabilities of current 3D grounding systems, creating one of the most significant challenges in embodied AI.
A common workaround to this scarcity constructs 3D feature fields from 2D features (e.g., CLIP embeddings) and performs text queries via dot products between the text and 3D embeddings. Although this provides good generalization, performance degrades with more detailed descriptions typical of real-world queries, as illustrated in Figure 3. From this perspective, the dual-encoder approach falls short of 3D grounding as it contradicts a core grounding requirement.
In this paper, we ask: can we combine the best part of both pipelines, i.e., training a powerful grounding model while still overcoming the data scarcity by learning from powerful 2D models? The key insight is that differentiable rendering provides a natural bridge between 3D and 2D. If we can predict 3D masks and render them into 2D views, we can supervise them using 2D foundation models that have been trained on internet-scale data. This approach could enable training 3D models without any 3D mask annotations.
We introduce Language-Indexed Field Transfer with Gaussian Splatting (LIFT-GS), which implements this idea as a practical training pipeline. Given a point cloud and language query, LIFT-GS predicts 3D Gaussian representations that can be rendered into multiple 2D views. These rendered masks are then supervised using pseudo-labels from 2D foundation models. We show results where SAM provides mask supervision (Kirillov et al., 2023), and CLIP or LLaMa provide language understanding (Radford et al., 2021; Meta AI, 2024). The approach effectively distills internet-scale 2D knowledge into 3D understanding.
This render-supervised formulation offers several key advantages. First, it is inherently architecture-agnostic; specifying only the outputs leaves flexibility in underlying model design. Second, this allows us to overcome fundamental scaling limitations by training a large transformer decoder instead of previous dual-encoder approaches (as shown in Fig 3(Zhu et al., 2023b; Gu et al., 2024). Third, the approach is highly practical: LIFT-GS operates directly on raw point clouds from sensors, such as the outputs from SLAM or SfM systems, eliminating the preprocessing and feature fusion required by methods like ConceptFusion (Jatavallabhula et al., 2023b; Arnaud et al., 2025), reducing inference time from  60 seconds to just 1 second.
Our experiments validate the effectiveness of this approach. LIFT-GS achieves state-of-the-art performance on standard 3D VLG benchmarks, with 25.7% mAP on open-vocabulary instance segmentation (vs. 20.2% previous SOTA) and consistent 10-30% relative improvements on referential grounding tasks. More importantly, we observe that across data scales for SFT, pretraining effectively ”multiplies” the fine-tuning dataset by approximately a constant factor (2x). That is, a pretrained model with 50% of fine-tuning data matches the performance of training from scratch with 100% data. This somewhat counterintuitive observation indeed matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021), and the fact that this scaling coefficient remains constant without diminishing returns across data scales adds an empirical data point that 3D VLG currently operates in the very low-data regime 111(Hernandez et al., 2021) define a “low-data regime as having 10% or less of the amount of data it would take to get to 99% of the performance that infinite data would yield.”.
The implications extend beyond 3D grounding. Render supervision is powerful, but we demonstrate that it can serve as a bridge for large-scale knowledge transfer from 2D foundation models to 3D models. Any 3D/4D task with renderable outputs can potentially leverage 2D supervision. As 2D foundation models continue to improve and scale, 3D models trained using render-supervised distillation are positioned to benefit. This opens the possibility of training 3D understanding models at the scale of 2D datasets–which would represent a fundamental shift from the current paradigm of limited 3D annotations.
To summarize, our contributions are:
  • A render-supervised training pipeline for 3D vision-language grounding that requires only 2D supervision. We show how differentiable rendering enables training 3D models with 2D losses, eliminating dependence on scarce 3D annotations.
  • Demonstrating a pseudo-labeling strategy for distilling 2D foundation models into 3D. LIFT-GS shows using SAM, CLIP, and LLMs to generate 2D supervision.
  • State-of-the-art performance in realistic evaluations. LIFT-GS achieves SOTA results using sensor point clouds common in embodied settings, with detailed ablations revealing scaling properties.
Figure 2: 3D Referential Grounding. For each mentioned instance in a text description, predict a 3D mask and map it to corresponding text tokens.
@@ 2 Related Work
2.1 The Data Scarcity Challenge in 3D VLG
3D Vision-Language Grounding (3D VLG) maps language descriptions to corresponding 3D masks or bounding boxes in observed scenes (Yuan et al., 2021; Roh et al., 2021; Yang et al., 2021). Despite its fundamental importance for embodied AI, existing 3D VLG datasets contain only thousands of annotated scenes (Dai et al., 2017; Yeshwanth et al., 2023) compared to billions of images used for training large multimodal models (Meta AI, 2024).
This scarcity stems from prohibitive annotation costs. Creating 3D instance masks requires minutes per example even with assisted tools (Dai et al., 2017), and human verification remains necessary (Arnaud et al., 2025; Majumdar et al., 2024). All existing 3D VLG methods require ground-truth 3D masks or bounding boxes during training (Yuan et al., 2021; Zhu et al., 2023c, 2024; Zhang et al., 2024), with many also requiring them during inference (Fang et al., 2024; Zhang et al., 2023b). This creates a fundamental scaling barrier–these approaches cannot leverage the vast 2D data that has driven progress in 2D understanding.
2.2 Bridging 2D and 3D: From Lifting to Learning
2D-to-3D Lifting via Optimization. Recent work addresses data scarcity by lifting 2D models to 3D through per-scene optimization. Methods use depth unprojection with heuristic merging (e.g., voxel voting) (Jatavallabhula et al., 2023a; Zhou et al., 2024; Xu et al., 2023b) or differentiable rendering to optimize 3D representations matching 2D features (Kerr et al., 2023; Kim et al., 2024; Gu et al., 2024) or masks (Cen et al., 2023; Xu et al., 2023a). While these leverage 2D foundation models, they suffer from: (1) slow optimization (minutes per scene), (2) accumulated errors from reconstruction and merging, and (3) inability to improve with more data. The fixed lifting pipelines may be a bottleneck that explains why LIFT-GS outperforms models trained on these lifted pseudolabels (Genova et al., 2021; Peng et al., 2023).
Render-Supervised Learning. Differentiable rendering enables training 3D models directly by rendering their predictions into 2D and supervising via 2D losses. While initially used for reconstruction (Hong et al., 2024; Tang et al., 2024; Cao et al., 2024; Szymanowicz et al., 2025),  (Irshad et al., 2024; Zhu et al., 2023b) use it to representation learning, like PonderV2 adds CLIP losses on rendered pixels (Zhu et al., 2023a). However, existing 3D VLG methods have significant limitations: PonderV2 only trains encoders and relies on ground-truth category labels, while other methods apply simple photometric losses without leveraging 2D VLMs. LIFT-GS overcomes these challenges by extending render-supervision in two key ways: (1) it enables training unified encoder-decoder architectures for 3D VLG tasks, and (2) it performs knowledge distillation from 2D foundation models (VLMs, SAM, CLIP) through pseudo-labeling, removing the need for 3D annotations during pretraining.
2.3 Architecture: From Dot-Products to Joint Attention
Figure 3: 3D grounding with CLIP-style (dual-decoder) method. Grounding heatmaps from a representative approach (Guo et al., 2024). Heatmaps are computed using dot product similarity between visual tokens and text tokens (as in the CLIP objective), encoded independently. This performs effectively with very short prompts, but fails with more detailed queries, as shown in the image. LIFT-GS addresses this by jointly predicting tokens using a transformer decoder with expressive attention masks (see Figure 6 and experiments).
Limitations of Dual-Encoder Methods. Almost all prior 3D grounding approaches that distill from 2D vision-language models eventually compute the mask as the dot-product similarity between 3D features and text embeddings (Radford et al., 2021; Guo et al., 2024; Qin et al., 2023; Kerr et al., 2023; Peng et al., 2023). When modality embeddings are computed independently, as is the case in CLIP, these “dual-encoder” models behave as bag-of-words systems (Yuksekgonul et al., 2022). Using these encoders causes 3D models to inherit these fundamental limitations; as shown in the limited ability of 3D dual-encoder models to handle relational language crucial for referential grounding (e.g., ”the chair next to the table”) (Fig. 3).
Multimodal Decoders. Transformer decoders address the bag-of-words behavior by jointly processing the modalities together through learned attention mechanisms that enable proper handling of spatial relationships and multi-object references. Following large multimodal language models (LLaMA 3, GPT-4o, and Qwen 2.5), as well as recent 3D VLG SotA (Kamath et al., 2021; Jain et al., 2025; Arnaud et al., 2025), LIFT-GS employs a decoder-based architecture. LIFT-GS introduces the grounding loss described in Sec. 3.2 to train the decoder.
2.4 Foundation Model Distillation at Scale
Recent work on scaling up pseudolabeling pipelines shows that although 2D foundation models are increasingly capable (Hong et al., 2023; Arnaud et al., 2025), they currently exhibit significant limitations in spatial understanding over multiple frames, frequently hallucinating 3D spatial relationships (Majumdar et al., 2024; Yang et al., 2024). This is why LIFT-GS uses pseudolabels for pretraining: just as LLMs require supervised fine-tuning (SFT) to align noisy internet text with desired behaviors, LIFT-GS leverages noisy pseudolabels for large-scale pretraining, then uses 3D VLG SFT for state-of-the-art performance.
Our scaling analysis reveals suggests that even imperfect 2D spatial understanding generates meaningful training signal (multiplying fine-tuning data effectiveness by 2× in our experiments). As 2D models advance in spatial reasoning, this transfer benefit should amplify, potentially reducing reliance on 3D annotations and moving toward more zero-shot spatial understanding.
This positions our approach at the intersection of three key trends: (1) the shift from optimization-based to learning-based 3D understanding, (2) the move from dual-encoder to multimodal decoder architectures for complex language grounding, and (3) the emergence of foundation model distillation as a solution to 3D data scarcity. Our experiments demonstrate that this approach not only achieves state-of-the-art performance but also exhibits strong scaling properties, suggesting significant potential for using cross-scene render-supervised distillation with improved multimodal foundation models and in other settings in besides 3D VLG.
@@ 3 Method
LIFT-GS is a (pre-)training pipeline for 3D VLG without 3D GT annotation. During training, it renders the predicted 3D masks from the target viewpoints for 2D supervision; during testing, 3D masks are used as outputs. This section covers the 3D VLG formalism in Sec.3.1, 2D loss used for supervision in Sec.3.2, pseudolabel generation in Sec.3.4, and implementation details in Sec.3.3.
@@ 3.1 Task Formulation
3.1.1 3D Vision-Language Grounding
Figure 2 shows an example using a point cloud input ($\mathbf{P})$ and the text query ($\mathbf{Q}$) “the black chair close to the table near the wall”. Following MDETR (Kamath et al., 2021) and (Jain et al., 2025), LIFT-GS outputs a set of ($m=256$) 3D mask candidates $\mathbf{M}$, and the correspondence matrix $\mathbf{C}$
@@
Matrix $\mathbf{C}\in\mathbb{R}^{m{\times}|Q|}$: indicates the correspondence (i.e. probability logits) between $m$ 3D mask candidates and $|Q|$ text tokens, enabling the mapping of each text query to its most probable mask candidate based on the logits.
3D Mask $\mathbf{M}\in\mathbb{R}^{m{\times}N}$: stores mask logits over $N$ Gaussian primitives for $m$ 3D mask candidates. Each logit $\mathbf{M}_{i,j}$ represents the probability (after applying sigmoid) that the $j$-th Gaussian primitive is included in the $i$-th mask candidate.
This dual design elegantly handles two key challenges: (1) text tokens can refer to multiple instances (e.g., “the chairs”), and (2) instances can be referenced multiple times with different descriptions throughout the text (e.g. both ”chair” and ”it”). It also extends readily to segmenting other modalities by adding more mappings.
The pointcloud has $|P|$ points and is of shape $|P|\times 6$ (XYZ + RGB), and the query token embeddings and are a matrix of size: $\mathbf{Q}\in\mathbb{R}^{|Q|{\times}F_{Q}}$. The text mapping $\mathbf{C}$ and Gaussian/point cloud mapping $\mathbf{M}$ are shown, with each instances highlighted in a different color in Figure 2.
3.1.2 3D VLG with Gaussian Masks $\mathbf{M}$
Existing 3D VLG methods are limited by costly point cloud annotation, where annotation for point cloud masks is by far the costliest and slowest step in 3D VLG data collection. This is because human annotators must carefully segment the mask using a brush or assisted tool, which takes on the order of minutes per example (Dai et al., 2017). Even with model assistance, human verification is required in practice (Arnaud et al., 2025; Majumdar et al., 2024). This restricts training to only thousands of scenes, making data scarcity the key bottleneck.
LIFT-GS addresses this by leveraging differentiable rendering, distilling knowledge from multimodal vision-language models trained on billions of images. It predicts the 3D Gaussians from the point cloud input feed-forwardly, which are used for rendering later. 3D Gaussians are represented with xyz locations $\mathbb{R}^{N\times 3}$, covariance matrices $\mathbb{R}^{N\times 6}$, color matrices $\mathbb{R}^{N\times 3}$ and feature embeddings $\mathbb{R}^{N\times F}$. LIFT-GS predicts $\mathbf{G}\in\mathbb{R}^{N\times(m+12+512)}$ containing $m$ masks plus 12 channels for shape/location/color and 512 for feature loss:
[…]
In practice, training the masks using differentiable rendering adds only a small overhead during training. With the shapes of the inputs and outputs specified, the sections below describe each component: the losses used, pseudolabel generation, and the model architecture.
3.2 Losses
With differentiable rendering, LIFG-GS enables training 3D VLG models using the simplest 2D grounding losses. During training, LIFT-GS only requires (sparse) point cloud, 2D posed images as inputs without any other 3D annotations, which data can be easily obtained from RGB-D videos or SfM (Yang et al., 2025; Wang et al., 2025a, 2023, b; Leroy et al., 2024) as done in  (Szymanowicz et al., 2025).
LIFT-GS utilizes two groups of losses to train the model: $\mathcal{L}_{\text{ground}}$ for grounding (Jain et al., 2025), and per-pixel losses $\mathcal{L}_{\text{PP}}$ commonly used to improve results with differentiable rendering. For reference, rendered 2D masks, rgb images, and feature maps are denoted as $\tilde{\mathbf{M}}_{\text{2D}}\in\mathbb{R}^{H\times W\times m}$, $\tilde{\mathbf{F}}_{\text{2D}}\in\mathbb{R}^{H\times W\times F}$, and $\tilde{I}\in\mathbb{R}^{H\times W\times 3}$, respectively.
Their corresponding ground-truth counterparts have analogous shapes: $\mathbf{M}_{\text{2D}}$, $\mathbf{F}_{\text{2D}}$, and $I$, where $\mathbf{M}_{\text{2D}}\in\mathbb{R}^{H\times W\times K}$.
3.2.1 Grounding losses:
LIFT-GS uses the MDETR-style mask grounding loss: $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{mask}}$ with $\mathbf{d}_{\text{match}}$ (matching distance). Since the number of predicted and ground-truth masks may differ, we apply Hungarian matching to pair them based on $\mathbf{d}_{\text{match}}$, and then optimize the matched predictions using $\mathcal{L}_{\text{ground}}$. We use the 2D variant of these losses under 2D rendering supervision, which can be easily replaced by their 3D counterpart (on 3D Gaussian centers) when 3D labels are available.
Cross Entropy Loss, $\mathcal{L}_{\text{CE}}$: this loss supervises the correspondence between 3D mask candidates and input language tokens (i.e., the matrix $\mathbf{C}$) by framing it as a (soft) classification problem. Recent work suggests that using BCE may produce sharper masks for long VLG text queries (Jain et al., 2025; Arnaud et al., 2025; Zhai et al., 2023), echoing similar findings in image segmentation (Cheng et al., 2021a).
Mask loss, $\mathcal{L}_{\text{mask}}$: this loss supervises the predicted 3D masks by comparing them to paired ground-truth 3D masks. Following SAM (Kirillov et al., 2023), we apply a combination of Focal (Lin et al., 2017) and Dice (Sudre et al., 2017) losses to supervise the predicted 3D masks effectively.
Optimal matching, $\mathbf{d}_{\text{match}}$: this function measures the pairwise distance between the 3D grounding results and the ground-truth values, and is used for matching. It is implemented similarly to $\mathcal{L}_{\text{ground}}$ but with different loss weights.
The model always predicts the maximum number of instances (256) but avoids false positive detections by matching the unused instance to a special no-match text token (Jain et al., 2021). The maximum of 256 was not a major limitation in practice, but could be increased.
3.2.2 Per-Pixel Losses
While grounding loss alone is sufficient for stable pretraining (Table 4), LIFT-GS benefits from joint training with additional photometric and feature losses for faster convergence. These losses are only used during pretraining, not for finetuning with 3D annotations.